# Debiasing Alternative Data for Credit Underwriting Using Causal Inference

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 5th ACM International Conference on AI in Finance; November 14–16, 2024; Brooklyn, NYbooktitle: Workshop on Explainable AI in Finance (XAIFIN2024), November 15, 2024, Brooklyn, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Causal reasoning and diagnosticsccs: Computing methodologies Probabilistic reasoningccs: Social and professional topics Computing / technology policyccs: Applied computing Law, social and behavioral sciences
@@ 2. Building a causal model of fairness
Block diagram of a black box machine learning model
@@
A three node CBN
@@
A five node CBN
@@ 3. Defining unlawful discrimination using causal models
A five node CBN
@@
A five node CBN
@@ 4. Incorporating alternative data into a causal model
A four node CBN
@@
A four node CBN
@@
An eight node CBN
@@
An eight node CBN
@@ 6.2. Results
Graph showing test results

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: 10.48550/arXiv.2410.22382conference: 5th ACM International Conference on AI in Finance; November 14–16, 2024; Brooklyn, NYbooktitle: Workshop on Explainable AI in Finance (XAIFIN2024), November 15, 2024, Brooklyn, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Causal reasoning and diagnosticsccs: Computing methodologies Probabilistic reasoningccs: Social and professional topics Computing / technology policyccs: Applied computing Law, social and behavioral sciences
@@ 2. Building a causal model of fairness
A black box ML model with data $X$ as input and decision $D$ as output. A protected attribute $A$ is also fed into the ML model. The decision $D$ leads to an outcome $Y$, which is then fed back into data $X$.
@@
A causal Bayesian network with protected attribute $A$ causing latent mediator $W$, which in turn causes outcome $Y$.
@@
A causal Bayesian network where data $X_{W}$ is a measure of mediator $W$, which causes decision $D$. This in turn causes outcome $Y$.
@@ 3. Defining unlawful discrimination using causal models
A red edge from the protected attribute $A$ to decision $D$ indicates bias due to overt discrimination. This also leads to a red edge from decision $D$ to outcome $Y$, also indicating bias.
@@
A causal Bayesian network where all of the nodes and edges are red to indicate bias due to covert discrimination.
@@ 4. Incorporating alternative data into a causal model
Demographic variable $Z$ partially explains away the causal effect between protected attribute $A$ and mediator $W$. The red edges indicate bias due to the spurious proxy effect that $Z$ has on $W$ through $A$.
@@
Proxy variable $P$ has no causal effect on mediator $W$. The red edges indicate bias due to the spurious proxy effect that $P$ has on $W$ through $A$.
@@
A red edge from alternative data $X_{Z}$ to decision $D$ indicates bias due to redlining and proxy discrimination.
@@
A blue edge from $A=a^{\prime}$ indicates a bias correction from the use of $X_{Z}$ on the decision $D$.
@@ 6.2. Results
The counterfactual model is able to recover most of the loss in ROC/AUC from the unawareness model.